OpenStudy (anonymous):

Will give medal What value of s makes the statement true? (–7 + s) + (–14) = (12 + (–7)) + (–14) A. –30 B. –12 C. –7 D. 12

OpenStudy (anonymous):

-7+s - 14 = 12 - 7 - 14 -21 + s = -9 s = 12

OpenStudy (anonymous):

does that make sense?